OpenStudy (anonymous):

Amistre64 I NEED your help!!!!!

OpenStudy (amistre64):

how can I help?

OpenStudy (anonymous):

Well I took a screen shot of this horrible problem I can't figure out! Imagine that!!! ; )

myininaya (myininaya):

i want to see the screen shot :)

OpenStudy (amistre64):

does it have a 2 in it lol

OpenStudy (amistre64):

they want you to do int by parts and put it in table format

OpenStudy (anonymous):

There it is! I cannot figure out how they did that addition in the end with those stupid e's and stuff! BTW I got a 100% on my calculus exam this morning! Thanks for the help!

OpenStudy (amistre64):

yay!!

OpenStudy (amistre64):

the table I set up is like this... let me make a pic

OpenStudy (anonymous):

Just so you know, two things bother me about this: the e^5x/125 and how they got the 75x^2-130x+26

OpenStudy (amistre64):

int 'v' will of course be the e^5x

OpenStudy (amistre64):

derive U will be 3x^2 -4x

OpenStudy (anonymous):

wait...what's int 'v'? Speak my language! You know, the language of imbeciles!

OpenStudy (amistre64):

integrate V

OpenStudy (amistre64):

this is my set up so far

OpenStudy (anonymous):

Oh, I see you mean the values for the table, right?

OpenStudy (amistre64):

does this make sense?

OpenStudy (amistre64):

i cant open your pdfs myin

OpenStudy (anonymous):

that most definitely does make sense! I get that part; it's the multplying through that I don't get. Can you help>

OpenStudy (anonymous):

I ouldn't open that attachment, either.

OpenStudy (amistre64):

now we simply multiply from right to left; which is why i offset it to make things easier

myininaya (myininaya):

:(

OpenStudy (anonymous):

tell me how you get 75x^2! I see an x^3 come out of that!!!!!

OpenStudy (amistre64):

after that its just simplifying if I did it right :)

myininaya (myininaya):

i guess i should save my files as jpegs right?

OpenStudy (amistre64):

jpegs come out quickly; but doesnt really matter

OpenStudy (amistre64):

i got what they got; but like I said; I stagger my table so I dont have to fool with the diagonal multiplication

OpenStudy (amistre64):

we can factor out a e^5x; and the bottom denoms are gonna have to be 125 to combine them; so \[\frac{e^{5x}}{125}\] is pulled out as a factor

OpenStudy (amistre64):

the rest is just adding up the fractions to simplify: \[\frac{3x^2-4x}{5}*\frac{25}{25}= \frac{75x^2 -100x}{125}\]

OpenStudy (amistre64):

\[\frac{6x-4}{25}*\frac{5}{5}=\frac{30x -20}{125}\]

OpenStudy (amistre64):

\[75x^2 -100x -(30x-20)\] \[75x^2 -100x -30x +20\] \[75x^2 -130x +20\]

OpenStudy (amistre64):

i forgot the last \(+6\) on that end lol

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

i am trying to figure out how they get those figures, So no, it doesn;t make sense, cuz I can't do the math.

OpenStudy (amistre64):

themath is just going back to adding and subtract fractions

OpenStudy (amistre64):

find a common denominator and redress the fractions to match

OpenStudy (amistre64):

\[{a\over 5}-{b\over 25}+{c\over 125}\] needs common denominators to simplify

OpenStudy (amistre64):

the common denom is: 125; so we redress the fractions so they are all wearing the same pants

OpenStudy (amistre64):

\[(\frac{a}{5}*{25\over25})-({b\over 25}*{5\over5})+{c\over 125}\]

OpenStudy (amistre64):

\[{25a \over 125} - {5b \over 125} +{6\over 125}\]

OpenStudy (amistre64):

pull out the common \({1\over 125}\) and fill in your a,b, and c

OpenStudy (amistre64):

well not the 'c' perse; thats just a 6 over there

OpenStudy (amistre64):

\[\frac{e^{5x}}{125}*(25(3x^2-4x)-5(6x-4)+6)\]

OpenStudy (amistre64):

do you see where they get it from yet?

OpenStudy (anonymous):

Trying to get a splinter out of my sons toe, please be patient.

OpenStudy (amistre64):

just cut the toe off :)

OpenStudy (anonymous):

tried that..,the butter knife just wasn't sharp enough!

OpenStudy (anonymous):

let me review your posts up there for a minute or two...the drama with 6yr olds just never ends!!!

OpenStudy (amistre64):

used ta get splinters all the time as a carpenter; just let them fester for abt a week and they pop out like a zit :)

OpenStudy (anonymous):

oh now that is just disgusting!!!! ; )

OpenStudy (amistre64):

lol... used to get these nasty one from yellow pine; they tips break off under the skin. first one I had I dint know it and it was sore for abt a week later; I finally got to meesing with it; just messaging the area and up poped the 'core' :)

OpenStudy (amistre64):

the ones i still cringe over are the ones that get tipped off under the fingernails..... shudder

OpenStudy (anonymous):

I still don't get that fraction thing.

OpenStudy (anonymous):

I am looking at it in all the different posts and just cannot pull it together to make sense/1 Please help me!

OpenStudy (amistre64):

simply put: how do we add: 3/4 + 2/3 ?

OpenStudy (anonymous):

you need a common denom of 12

OpenStudy (amistre64):

right; same step in your problem; except the common is 125

OpenStudy (amistre64):

now how do we redress the fractions to get a 12 under them?

OpenStudy (anonymous):

OMGosh! you're amazing cuz I GET IT!!!

OpenStudy (amistre64):

:) good

OpenStudy (anonymous):

let me try the math to make sure I get the same answer they have and then I will come back in a sec. Is that ok?

OpenStudy (amistre64):

doors are closing in about 10 mins ... then im off to the house for the night

OpenStudy (anonymous):

I'll bet your kids are really really smart, right?

OpenStudy (amistre64):

lol ..... they are ..umm.... can be :)

OpenStudy (anonymous):

oh, you mean you'll be gone?

OpenStudy (amistre64):

they wont let me stay in the library once the lock it up lol .... and my veranda is to far to travel to after coming to the college

OpenStudy (anonymous):

one more thing...with the signs for these? the positive negative posiotive? Is that always the same? The sign of the first one is + then - then +

OpenStudy (amistre64):

its alternating yes; +-+-+-+-+-+-+ ..... start with + and cycle it

OpenStudy (amistre64):

thats a sideeffect of doing 'multiple' integration by parts

OpenStudy (amistre64):

and the table is just a conveint way to organize 'multiple' integration by parts

OpenStudy (anonymous):

Thank you still trying to work it, but I get the picture now. You're so kind and patient; I really don't know what I would have done here in calc without your input. My A is due to partly your great help!

OpenStudy (amistre64):

:) thanx

OpenStudy (anonymous):

Talk to you maybe tomorrow?

OpenStudy (amistre64):

tomorrow should be an alnighter for me; college is closed so I go to the place with my veranda

OpenStudy (anonymous):

Oh, fun. I remember those all-nighters in college! Course I wasn't 40 back then and could do it! : )

OpenStudy (amistre64):

:) Gnite

OpenStudy (anonymous):

nite.